# AC3D: Analyzing and Improving 3D Camera Control in Video Diffusion Transformers

### v2

@@ Abstract
Numerous works have recently integrated 3D camera control into foundational text-to-video models, but the resulting camera control is often imprecise, and video generation quality suffers. In this work, we analyze camera motion from a first principles perspective, uncovering insights that enable precise 3D camera manipulation without compromising synthesis quality. First, we determine that motion induced by camera movements in videos is low-frequency in nature. This motivates us to adjust train and test pose conditioning schedules, accelerating training convergence while improving visual and motion quality. Then, by probing the representations of an unconditional video diffusion transformer, we observe that they implicitly perform camera pose estimation under the hood, and only a sub-portion of their layers contain the camera information. This suggested us to limit the injection of camera conditioning to a subset of the architecture to prevent interference with other video features, leading to 4×4$\times$4 × reduction of training parameters, improved training speed and 10% higher visual quality. Finally, we complement the typical dataset for camera control learning with a curated dataset of $20$K diverse dynamic videos with stationary cameras. This helps the model disambiguate the difference between camera and scene motion, and improves the dynamics of generated pose-conditioned videos. We compound these findings to design the Advanced 3D Camera Control (AC3D) architecture, the new state-of-the-art model for generative video modeling with camera control.
Figure 1: Camera-controlled video generation. Our method enables precise camera controllability in pre-trained video diffusion transformers, allowing joint conditioning of text and camera sequences. We synthesize the same scene with two different camera trajectories as input. The inset images visualize the cameras for the videos in the corresponding columns. The left camera sequence consists of a rotation to the right, while the right camera visualizes a zoom-in, up, and zoom-out trajectory.
@@ 1 Introduction
Foundational video diffusion models (VDMs) trained on internet-scale data, acquire abundant knowledge about the physical world [10]. They not only learn appearance and plausible 2D dynamics, but they also have abundant understanding of 3D structure [7]. However, most of this knowledge is stored implicitly within the model, as these models do not expose fine-grained control mechanisms, such as camera motion control.  We recently witnessed a surge of works that bring 3D camera control into foundational video models [149, 39, 165], but the control they provide is not very precise, and the synthesis quality is often compromised [6]. We analyze camera motion control in video diffusion models from first principles, and develop several findings that allow us to incorporate precise 3D camera conditioning without degrading synthesis quality.  To perform our analysis we train a 11.5B-parameters VDiT (video latent diffusion transformer) [100] on a dataset of 100M text/video pairs. On this model, we perform three key studies. With what we learn, we adapt the camera control solution from VD3D [6] from a pixel-based to latent-based diffusion model, and significantly increase its performance.
@@ 2 Related work
Our approach builds on recent advancements in 2D video diffusion models. One prominent technique in this area enhances text-to-image models by adding temporal layers to support video generation [8, 119, 151, 37, 7, 37]. While these methods use the U-Net architecture, more recent ones [10, 92, 93, 167] have been adapting transformer-based architectures for more scalable, realistic, and highly dynamic scene generation. We are interested in controlling the camera movements during the generation process of recent transformer-based video models based on precise camera extrinsics, i.e., cameras represented as rotation and translation sequences for each frame.
@@
Early 4D generation works [2, 163] used 4D GANs to learn category-specific generators with an underlying dynamic 3D representation. More recent approaches  [120, 5, 194, 81, 5, 4] have tackled 4D generation by distilling motion priors from pre-trained video diffusion models into an explicit 4D representation, enabling category-agnostic 4D generation. Follow-up works investigate image or video conditioned 4D generation  [109, 191, 171, 98, 194, 81, 33, 180, 76] instead of pure text inputs, improving flexibility in the generation process. While most of these works are object-centric, recent approaches [4, 158] shifted towards more complex scenes, including methods [23, 173] which model the background. However, all these methods are optimization-based, i.e., each scene is generated independently from scratch. Recently, L4GM [110] proposes a feed-forward 4D generator trained on object-centric synthetic 4D data. While these approaches are explicit and provide space-time control, they are limited in their photorealism compared to recent 2D video diffusion models. We investigate dynamic 3D scene generation from a different perspective by extending pre-trained video diffusion models with 3D camera control.
@@
Concurrent approaches [192, 157, 193, 183, 175, 68] further improve camera control in U-Net-based architectures, while another work [22] tackles video diffusion transformer. However, the scene and visual quality is still limited in that approach. DimensionX [127] controls space and time in video diffusion transformers but the camera trajectories are pre-defined and not continuous. [51] explores pose estimation with a video DiT by pairing it with DUSt3R [145] and fine-tuning, while we perform linear probing without any training to assess its existing camera knowledge.
@@ 3 Method
We first describe our base video diffusion model (Sec. 3.1), and the baseline camera control method built on top of it (Sec. 3.2). Then, we proceed with the analysis of motion (Sec. 3.3), linear probing (Sec. 3.4) and dataset biases (Sec. 3.5), and additional insights on how to build an effective camera control (Sec. 3.6)
@@ 3.1 Base model (VDiT)
Following Sora [10], most modern foundational video generators use the diffusion framework [122, 45] to train a large-scale transformer [139] in the latent space of a variational autoencoder [64, 111] to generate videos from text descriptions. We adopt the same design and, for a base video model, pre-train an 11.5B-parameters Video DiT model [100] with 32-blocks of 4,096 hidden dimensionality for text-to-video generation in the latent space of CogVideoX [167] $16$-channels $4{\times}8{\times}8$-compression autoencoder with the rectified flow diffusion parametrization [85]. We use the T5 [108] encoder to produce text embeddings, which are passed into VDiT via cross-attention. We train our base model on a large-scale dataset of images and videos with text annotations, with resolutions ranging from $17{\times}144{\times}256$ to $121{\times}576{\times}1024$. Such design is fairly standard and followed by many existing works with little deviation [102, 167, 32, 195], and we describe our particular architectural and training setups in detail in Appendix C.
@@ 3.2 VDiT with Camera Control (VDiT-CC)
To construct a baseline architecture for camera control, we implement ControlNet [186, 18] conditioning on top VDiT. Similarly to [149, 39, 6], we use the RealEstate10k [197] dataset, consisting of 65k (text, video, camera trajectory) triplets ${(\bm{x}_{n},\bm{y}_{n},\bm{c}_{n})}_{n=1}^{N}$ and train a new set of model parameters to input the camera information into the model. Camera trajectories $\bm{c}\in\mathbb{R}^{f\times 25}$ are provided in the form of camera extrinsics $\bm{C}_{f}\in\mathbb{R}^{4\times 4}$ and intrinsics $\bm{K}_{f}\in\mathbb{R}^{3\times 3}$ for each $f$-th frame $\bm{x}_{f}$.
@@
For base camera control, we adapt VD3D [6] since it was designed for transformer-based models and suits our setup the most, while other methods are built on top of UNet-based [112] backbones. We use Plücker camera representations [121, 16, 59, 6, 39], which are projected to the same dimensionality and resolution as the video tokens via a fully-convolutional encoder to produce camera tokens.  These camera tokens are processed by a sequence of lightweight DiT-XS blocks with 128 hidden dimensionality and 4 attention heads each. To mix the camera information with the video tokens of VDiT, we found a summation to be sufficient, which we perform before each main DiT block. We also found it useful to perform cross-attention from video tokens to camera tokens as a form of a feedback connection [71]. We illustrate this architecture VDiT-CC in Figure 2; see implementation details in Appendix C.
@@
Keeping the VDiT backbone frozen, we train the new parameters with a rectified flow objective [85] and standard (location of 0 and scale of 1) logit-normal noise distribution [28]. Similar to prior works[6, 150], we apply a $10\%$ camera dropout to support classifier-free guidance (CFG) [44] later. Notably, we train VDiT-CC only at the $256^{2}$ resolution: since camera motion is a low-frequency type of signal (which is fully unveiled at lower resolutions) and the main VDiT backbone is frozen, we found that such design generalizes to higher resolutions out-of-the-box. During inference, we input text prompts and camera embeddings with classifier-free guidance at each time step.
Figure 3: Comparing motion spectral volumes for scenes with different motion types. Videos with camera motion (purple) exhibit stronger overall motion than the videos with scene motion (orange), especially for the low-frequency range, suggesting that the motion induced by camera transitions is heavily biased towards low-frequency components.
@@
This baseline model, being built on top of our powerful VDiT, already achieves decent-quality camera control. However, it struggles with degraded visual quality, reduced scene motion, and sometime the camera control gets ignored. To improve the design, we analyze our VDiT backbone to understand how camera motion is modeled and represented by it, and then inspect VDiT-CC’s failure cases and where they arise to rigorously address them.
@@ 3.3 How is camera motion modeled by diffusion?
We start by analyzing how camera motion is modeled by a pre-trained video diffusion model (i.e., before camera control is incorporated). We hypothesize that the motion induced by changes in camera pose is a low-frequency type of signal and investigate the motion spectral volumes [75] of the generated videos at different steps of the denoising process. To perform this analysis, we generate $200$ diverse videos with our VDiT model with 80-steps denoising and manually annotate them into four categories: videos with only scene motion, videos with only camera motion, videos with both scene and camera motion, and others; see Appendix D for details. During generation, we save the denoised predictions at each denoising step and estimate optical flow to compute the motion spectral volumes.
@@
We visualize these lines with 95% confidence intervals in Figure 3. Videos with camera motion exhibit higher amplitudes than scene-motion-only videos for low-frequency components while having similar characteristics for high-frequency ones. This supports the conjecture that the camera motion is a low-frequency type of signal. We also depict an example of a generated video with both scene and camera motion with four denoising steps on Fig. 4(a): one can observe that the camera movement has been fully produced by $t{=}0.9$ (first $10\%$ of the rectified flow denoising process). In contrast, scene motion details like the hand movements of the subjects are not finalized even till $t{=}0.5$.
Inspired by this finding, we pose the question: when does exactly a video diffusion model determine the camera? To answer it, we plot aggregated spectral volumes for different timesteps in Figure 4(b), and also their ratio with respect to the last timestep $t{=}0$ (for which all the motion has been produced) We then inspect when different types of motion kick in during the denoising process. Figure 4(b) (right) shows that the low-frequency motion components fill up to ${\approx}{84}$% at $t{=}0.9$ (the first 10% of the denoising process), while high-frequency ones are not well complete until $t{=}0.6$.
An immediate consequence of this observation is that trying to control the camera later in the denoising trajectory is simply unnecessary and will not influence the manipulation result. In this way, instead of using the standard logit-normal noise level distribution of SD3 [28] with a location of 0.0 and scale of 1.0 (which we use by default for VDiT), we switch to using truncated normal with a location of 0.8 and scale of 0.075 on the [0.6, 1] interval to cover the early steps of the denoising rectified flow trajectory. Besides, we apply camera conditioning at inference time on the same [0.6, 1] interval. But surprisingly, we observe that it is unnecessary but also detrimental to the scene motion and overall visual quality. Following this insight, we restrict both our train-time noise levels and test-time camera conditioning schedules to cover only the first 40% of the reverse diffusion trajectory. As Sec. 4.3 shows, it improves FID and FVD by $14\%$ on average, and camera following by $30\%$ on MSR-VTT (the dataset which measures generalization to diverse, out-of-fine-tuning-distribution scenes). Besides, it enhances the overall scene motion, which we demonstrate qualitatively.
@@ 3.4 What does VDiT know about camera pose?
Foundational video models acquire rich knowledge about the physical world, and we hypothesize that they store information about the camera signal within their representations. To investigate this, we perform linear probing of our base VDiT model on the RealEstate10k [197] dataset (not seen during training) for camera extrinsics.
Specifically, we take 1,000 random 49-frame videos from RealEstate10K, feed them into VDiT under 8 noise levels $(1/8,2/8,...,1)$, and extract the activations for all 32 DiT blocks. Next, we split them into 900 train and 100 test videos and train a linear ridge regression model to predict the rotation pitch/yaw/roll angles and translation vectors for the entire viewpoint trajectory ($49\times 6$ target values in total). This results in $8\times 32$ trained models, and we report the rotation and (normalized) translation errors [39] on a held-out test set of 100 videos in Figure 5. Surprisingly, VDiT can accurately predict the camera signal, achieving minimum test errors of ${\approx}{0.025}$ for rotation and for ${\approx}{0.48}$ translation prediction. The knowledge quality increases around layer #9 and peaks in the range of #13-21. We reason that since the camera information in block #13 is stored in such a disentangled manner, then the model is using it to build other representations, and conditioning the camera at this point is risky and unnecessary and would interfere with other visual features. In this way, we propose to input the camera only in the first #8 blocks and leave the rest 24 DiT blocks unconditioned. We empirically found in Section 4.3, that this not only reduces the amount of trainable parameters ${\approx}{4}$ times and improves training speed ${\approx}{15}$%, but also enhances the visual quality ${\approx}{10}$%.

### v4

@@ Abstract
Numerous works have recently integrated 3D camera control into foundational text-to-video models, but the resulting camera control is often imprecise, and video generation quality suffers. In this work, we analyze camera motion from a first principles perspective, uncovering insights that enable precise 3D camera manipulation without compromising synthesis quality. First, we determine that motion induced by camera movements in videos is low-frequency in nature. This motivates us to adjust train and test pose conditioning schedules, accelerating training convergence while improving visual and motion quality. Then, by probing the representations of an unconditional video diffusion transformer, we observe that they implicitly perform camera pose estimation under the hood, and only a sub-portion of their layers contain the camera information. This suggested us to limit the injection of camera conditioning to a subset of the architecture to prevent interference with other video features, leading to a 4×4$\times$4 × reduction of training parameters, improved training speed, and 10% higher visual quality. Finally, we complement the typical dataset for camera control learning with a curated dataset of $20$K diverse, dynamic videos with stationary cameras. This helps the model distinguish between camera and scene motion and improves the dynamics of generated pose-conditioned videos. We compound these findings to design the Advanced 3D Camera Control (AC3D) architecture, the new state-of-the-art model for generative video modeling with camera control.
Figure 1: Camera-controlled video generation. Our method enables precise camera controllability in pre-trained video diffusion transformers, allowing joint conditioning of text and camera sequences. We synthesize the same scene with two different camera trajectories as input. The inset images visualize the cameras for the videos in the corresponding columns. The left camera sequence consists of a rotation to the right, while the right camera visualizes a zoom-out and up trajectory.
@@ 1 Introduction
Foundational video diffusion models (VDMs) trained on internet-scale data, acquire abundant knowledge about the physical world [12]. They not only learn appearance and plausible 2D dynamics, but they also have abundant understanding of 3D structure [9]. However, most of this knowledge is stored implicitly within the model, as these models do not expose fine-grained control mechanisms, such as camera motion control.  We recently witnessed a surge of works that bring 3D camera control into foundational video models [151, 41, 168], but the control they provide is not very precise, and the synthesis quality is often compromised [6]. We analyze camera motion control in video diffusion models from first principles, and develop several findings that allow us to incorporate precise 3D camera conditioning without degrading synthesis quality.  To perform our analysis we train a 11.5B-parameter VDiT (video latent diffusion transformer) [102] on a dataset of 100M text/video pairs. On this model, we perform three key studies. With what we learn, we adapt the camera control solution from VD3D [6] from a pixel-based to latent-based diffusion model, and significantly improve its performance.
@@ 2 Related work
Our approach builds on recent advancements in 2D video diffusion models. One prominent technique in this area enhances text-to-image models by adding temporal layers to support video generation [10, 121, 153, 39, 9]. While these methods use the U-Net architecture, more recent ones [12, 94, 95, 170] have been adapting transformer-based architectures for more scalable, realistic, and highly dynamic scene generation. We are interested in controlling the camera movements during the generation process of recent transformer-based video models based on precise camera extrinsics, i.e., cameras represented as rotation and translation sequences for each frame.
@@
Early 4D generation works [2, 166] used 4D GANs to learn category-specific generators with an underlying dynamic 3D representation. More recent approaches  [122, 5, 198, 83, 4] have tackled 4D generation by distilling motion priors from pre-trained video diffusion models into an explicit 4D representation, enabling category-agnostic 4D generation. Follow-up works investigate image or video conditioned 4D generation  [111, 195, 175, 100, 198, 83, 35, 184, 78] instead of pure text inputs, improving flexibility in the generation process. While most of these works are object-centric, recent approaches [4, 161] shifted towards more complex scenes, including methods [25, 177] which model the background. However, all these methods are optimization-based, i.e., each scene is generated independently from scratch. Recently, L4GM [112] proposes a feed-forward 4D generator trained on object-centric synthetic 4D data. While these approaches are explicit and provide space-time control, they are limited in their photorealism compared to recent 2D video diffusion models. We investigate dynamic 3D scene generation from a different perspective by extending pre-trained video diffusion models with 3D camera control.
@@
Concurrent approaches [196, 160, 197, 187, 179, 70] further improve camera control in U-Net-based architectures, while another work [24] tackles video diffusion transformer. However, the scene and visual quality is still limited in that approach. DimensionX [129] controls space and time in video diffusion transformers but the camera trajectories are pre-defined and not continuous. Chun-Hao et al. [53] explore pose estimation with a video DiT by pairing it with DUSt3R [147] and fine-tuning, while we perform linear probing without any training to assess its existing camera knowledge. CAT4D [154] proposes a multi-view video diffusion model fine-tuned from a multi-view model. By scaling multi-view synthetic data generation, SynCamMaster [8] and ReCamMaster [7] achieve impressive results on camera-controlled video generation with transformers.
@@ 3 Method
We first describe our base video diffusion model (Sec. 3.1), and the baseline camera control method built on top of it (Sec. 3.2). Then, we proceed with the analysis of motion (Sec. 3.3), linear probing (Sec. 3.4) and dataset biases (Sec. 3.5), and additional insights on how to build an effective model for camera control (Sec. 3.6)
@@ 3.1 Base model (VDiT)
Following Sora [12], most modern foundational text-to-video generators use the diffusion framework [124, 47] to train a large-scale transformer [141] in the latent space of a variational autoencoder [66, 113]. We adopt the same design and, for a base video model, pre-train an 11.5B-parameter Video DiT model [102] with 32 blocks of hidden dimension 4,096 for text-to-video generation. We use the rectified flow diffusion parametrization [87] and learn in the latent space of CogVideoX [170] (using an autoencoder with a $16$-channel output and compression factors of $4{\times}8{\times}8$ in the temporal and spatial dimensions). The T5 [110] encoder produces text embeddings, which are passed into VDiT via cross-attention. We train our base model on a large-scale dataset of images and videos with text annotations, with resolutions ranging from $17{\times}144{\times}256$ to $121{\times}576{\times}1024$. This design is fairly standard and followed by many existing works with little deviation [104, 170, 34, 199]; we describe our specific architectural and training setup in detail in Appendix D.
@@ 3.2 VDiT with Camera Control (VDiT-CC)
To construct a baseline architecture for camera control, we implement ControlNet [190, 20] conditioning on top of the VDiT. Similar to previous work [151, 41, 6], we use the RealEstate10k [201] dataset, consisting of 65k (text, video, camera trajectory) triplets ${(\bm{t}_{n},\bm{x}_{n},\bm{c}_{n})}_{n=1}^{N}$ and train a new set of model parameters to input the camera information into the model. Camera trajectories $\bm{c}\in\mathbb{R}^{f\times 25}$ are provided in the form of camera extrinsics $\bm{C}_{f}\in\mathbb{R}^{4\times 4}$ and intrinsics $\bm{K}_{f}\in\mathbb{R}^{3\times 3}$ for each $f$-th frame $\bm{x}_{f}$.
@@
For base camera control, we adapt VD3D [6] since it was designed for transformer-based models and suits our setup the most, while other methods are built on top of UNet-based [114] backbones. We use Plücker camera representations [123, 18, 61, 6, 41], which are projected to the same dimensionality and resolution as the video tokens via a fully-convolutional encoder to produce camera tokens.  These camera tokens are processed by a sequence of lightweight DiT-XS blocks with hidden dimension 128 and four attention heads each. To mix the camera information with the video tokens of VDiT, we use summation before each main DiT block. We also found it useful to perform cross-attention from video tokens to camera tokens as a form of a feedback connection [73]. We illustrate this model architecture, which we call VDiT-CC, in Figure 2; see implementation details in Appendix D. VDiT-CC describes the camera-controlled video model architecture used by AC3D, while AC3D describes our proposed work including analysis and additional adjustments based on the analysis.
@@
Keeping the VDiT backbone frozen, we train the new parameters with a rectified flow objective [87] and standard (location of 0 and scale of 1) logit-normal noise distribution [30]. Similar to prior works [6, 152], we apply a $10\%$ camera dropout to support classifier-free guidance (CFG) [46] later. Notably, we train VDiT-CC only at the $256^{2}$ resolution: since camera motion is a low-frequency type of signal (which can be observed at lower resolutions) and the main VDiT backbone is frozen, we found that our design generalizes to higher resolutions out-of-the-box. During inference, we input text prompts and camera embeddings with classifier-free guidance at each time step.
Figure 3: Average magnitude of motion spectral volumes along spatial, temporal offset, and video batch dimensions for scenes with different motion types. We compute the flow of each video in a sliding window manner with temporal offsets and average the frequencies across all offsets. Videos with camera motion (purple) exhibit stronger overall motion than the videos with scene motion (orange), especially for the low-frequency range, suggesting that the motion induced by camera transitions is heavily biased towards low-frequency components. Frequency refers to the temporal frequency.
@@
This baseline model, being built on top of our powerful VDiT, already achieves decent-quality camera control. However, it struggles with degraded visual quality and reduced scene motion, and sometimes, the camera control inputs are ignored. To improve the design, we analyze our VDiT backbone to understand how camera motion is modeled and represented. Then, we inspect VDiT-CC’s failure cases and where they arise to address them.
@@ 3.3 How is camera motion modeled by diffusion?
We start by analyzing how camera motion is modeled by a pre-trained video diffusion model (i.e., before camera control is incorporated). We hypothesize that the motion induced by changes in camera pose is a low-frequency type of signal and investigate the motion spectral volumes [77] of the generated videos at different steps of the denoising process. To perform this analysis, we generate $200$ diverse videos with our VDiT model with 80 denoising steps and manually annotate them into four categories: videos with only scene motion, videos with only camera motion, videos with both scene and camera motion, and others; see Appendix E for details. During generation, we save the denoised predictions at each denoising step and estimate optical flow to compute the motion spectral volumes.
@@
We visualize motion spectral volumes with 95% confidence intervals in Figure 3. Videos with camera motion exhibit higher amplitudes than scene-motion-only videos for low-frequency components while having similar characteristics for high-frequency ones. This supports the conjecture that the camera motion is a low-frequency type of signal. We also depict an example of a generated video with both scene and camera motion with four denoising steps on Fig. 4(a): one can observe that the camera movement has been fully produced by $t{=}0.9$ (first $10\%$ of the rectified flow denoising process). In contrast, scene motion details like the hand movements of the subjects are not finalized even till $t{=}0.5$.
Inspired by this finding, we pose the question: when exactly does a video diffusion model determine the camera pose? To answer this question, we plot aggregated spectral volumes for different timesteps in Figure 4(b). We also show the ratio with respect to the last timestep $t{=}0$ (i.e., when all motion has been generated). We then inspect when different types of motion appear during the denoising process. Figure 4(b) (right) shows that the low-frequency motion components fill up to ${\approx}{84}$% at $t{=}0.9$ (the first 10% of the denoising process), while high-frequency components are not well-modeled until $t{=}0.6$.
An immediate consequence of this observation is that trying to control the camera later in the denoising trajectory is simply unnecessary and will not influence the manipulation result. In this way, instead of using the standard logit-normal noise level distribution of SD3 [30] with a location of 0.0 and scale of 1.0 (which we use by default for VDiT), we switch to using truncated normal with a location of 0.8 and scale of 0.075 on the [0.6, 1] interval to cover the early steps of the denoising rectified flow trajectory. At inference time, we apply camera conditioning on the same [0.6, 1] interval. Surprisingly, we observe that not using truncation is detrimental to the scene motion and overall visual quality. Following this insight, we restrict both our train-time noise levels and test-time camera conditioning schedules to cover only the first 40% of the reverse diffusion trajectory. As Sec. 4.3 shows, this improves FID and FVD by $14\%$ on average, and camera following by $30\%$ on MSR-VTT (the dataset used to measure generalization to diverse, out-of-fine-tuning-distribution scenes). Further, truncated noise sampling enhances the overall scene motion.
@@ 3.4 What does VDiT know about camera pose?
Foundational video models acquire rich knowledge about the physical world, and we hypothesize that they store information about the camera pose within their representations. To investigate this, we perform linear probing of our base VDiT model on the RealEstate10k [201] dataset (not seen during training) for camera extrinsics.
Specifically, we take 1,000 random 49-frame videos from RealEstate10K, feed them into VDiT under 8 noise levels $(1/8,2/8,...,1)$, and extract the activations for all 32 DiT blocks. Next, we split the random videos into 900 train and 100 test videos and train a linear ridge regression model to predict the rotation pitch/yaw/roll angles and translation vectors for the entire viewpoint trajectory ($49\times 6$ target values in total). This results in $8\times 32$ trained models, and we report the rotation and (normalized) translation errors [41] on a held-out test set of 100 videos in Figure 5. Surprisingly, VDiT can accurately predict the camera pose, achieving minimum test errors of ${\approx}{0.025}$ for rotation and for ${\approx}{0.48}$ translation prediction. The knowledge quality increases around layer #9 and peaks in the range of #13–21. We reason that since the camera information in block #13 is stored in such a disentangled manner, then the model is using it to build other representations; hence, conditioning the camera in this block is risky and unnecessary and would interfere with other visual features, as shown in our ablations. In this way, we propose to input the camera conditioning only in the first #8 blocks and leave the remaining 24 DiT blocks unconditioned. We find in Section 4.3 that this not only reduces the number of trainable parameters by ${\approx}{4}$ times and improves training speed by ${\approx}{15}$%, but also enhances the visual quality by ${\approx}{10}$%.